OpenStudy (calculusxy):

It is a Saturday morning and Jeremy has discovered he has a leak coming from the water heater in his attic. Since plumbers charge extra to come out on weekends, Jeremy is planning to use buckets to catch the dripping water. He places a bucket under the drip and steps outside to walk the dog. In half an hour the bucket is 1/5 of the way full. a. What is the rate at which the water is leaking per hour? b. What is the longest that Jeremy can be away from the house before the bucket will overflow?

OpenStudy (calculusxy):

@agent0smith

OpenStudy (calculusxy):

@StudyGurl14

OpenStudy (calculusxy):

@triciaal

OpenStudy (calculusxy):

For A I got 1/5 every1/2 hour. For B I got 2 1/2 hours.

OpenStudy (triciaal):

In half an hour the bucket is 1/5 of the way full. let the full bucket = x 1/2*1hr = 1/5 *X on the left to get 1 hr multiply by 2 since equation do the same to right hand side

OpenStudy (triciaal):

a. What is the rate at which the water is leaking per hour? per hour means for every one hour

OpenStudy (triciaal):

b. What is the longest that Jeremy can be away from the house before the bucket will overflow? this is asking to find the value of X, the full bucket

OpenStudy (calculusxy):

2/5 per hour

OpenStudy (calculusxy):

For part B I am incorrect?

OpenStudy (triciaal):

yes 2/5 of X per hour

OpenStudy (triciaal):

b) (2/5) X = 1 hr solve for X

OpenStudy (triciaal):

did you get X = 5/2 = 2 1/2 hrs ?

OpenStudy (calculusxy):

Yes.

OpenStudy (triciaal):

great

OpenStudy (calculusxy):

I have another question.

OpenStudy (calculusxy):

@triciaal

OpenStudy (triciaal):

open a new post